OpenStudy (anonymous):

What is the value of the expression 7 + 52 × (7 – 3)? a. 74 b. 107 c. 128 d. 179

OpenStudy (anonymous):

do you know PEDMAS or PEMDAS

OpenStudy (anonymous):

yeaah

OpenStudy (anonymous):

so do parenthesis first

OpenStudy (anonymous):

7 + 52 × (7 – 3) 7 + 52 x (4)

OpenStudy (anonymous):

i know but idk what to do after the parenthesis :-/

OpenStudy (anonymous):

is that what you are looking for?

OpenStudy (anonymous):

then multiplication

OpenStudy (anonymous):

7 + 208

OpenStudy (anonymous):

if x means multiplication, then I'm afraid neither of the answers are correct

OpenStudy (anonymous):

i got 215!

OpenStudy (anonymous):

well oh, it's 5 to the power of 2 not 52

OpenStudy (anonymous):

oh then it's a different story

OpenStudy (anonymous):

yup, that's why i got stuck

OpenStudy (anonymous):

okay, I guess you can do your own then

OpenStudy (mathstudent55):

\(7 + 5^2 \times (7 – 3)\) Parentheses first: \(7 + 5^2 \times \color{red}{(7 – 3)}\) \(7 + 5^2 \times \color{red}{4}\) Exponent next: \(= 7 + \color{green}{5^2} \times 4 \) \(=7 + \color{green}{25} \times 4\) Multiplication next: \(=7 + \color{blue}{25 \times 4}\) \(=7 + \color{blue}{100} \) Finally addition: \(=7 + 100 \) \(= 107\)